Now assume that the particle "chooses" a different way of traveling. Calculate the total amount of work W_ADC done by the electrostatic force on the charged particle as it moves from A to D to C.
Answer #1

\WABC = Fdcos? = (qE)(AC)cos? =(qE)(L/cos?)cos? = qEL
or WABC = WAB +WBC=qEL


qEL
conclusion: Coulumb forces are conservative
Answer #2

work done by the filed = change in potential enegy of the particle due to the field


a pic is need to lnow the configuratiopn of the points


so please provide a pic or some more info on the question
Answer #3
A to D to C figure ??